OpenStudy (anonymous):

find the exact value Given that cos theta = 15/17 with theta quadrant 1 , find sin 2theta and cos 2theta Plzz help

OpenStudy (solomonzelman):

 knowing SOHCAHTOA find the rest.

OpenStudy (solomonzelman):

@shorouq91, you can do it.

OpenStudy (anonymous):

If cosT=15/17 and T is in first quadrant then sinT >0 and sin^2T+cos^2T=1 so sin T=8/17. Thus from the fact that sin2a=2sinacosa and cos2a=1-2sin^2a you write : \[\sin2T=2*cosT*sinT=2*15/17*8/17=2*40/289=80/289\] \[\cos2T=1-2*\sin^2T=1-2*(5/17)^2=1-50/289=239/289\]

OpenStudy (anonymous):

im trying but i cant :(

OpenStudy (anonymous):

Hope it helps ;]

OpenStudy (anonymous):

yes it is thank you so much @Lolavitz

OpenStudy (anonymous):

who did u get 8

OpenStudy (anonymous):

Holy shiet, i forgot it is 15 not 5. So it is 2*15*8/17^2=250/289 My bad, thanks for noticing

OpenStudy (anonymous):

i'm confused now

OpenStudy (anonymous):

I made a mistake in calculations in multiplying. But the 8/17 i got from the fact that \[\cos^2T+\sin^2T=1\] which is real for every real number T. Thus we can write \[\sin^2T=1-\cos^2T+\] thus \[\sin^2T=1-225/289=289/289-225/289=(289-225)/289=64/289\] Thus sinT equals either -8/15 or 8/15 and since we know that T is in first quadrant so sinT >0 so sinT=8/15.

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{15}{17}\implies \cfrac{\textit{adjacent side}}{hypotenuse}\implies \cfrac{a}{c}\qquad thus\\ \quad \\ a = 15\qquad c= 17\qquad b=\square ?\qquad \textit{using pythagorean theorem}\\ \quad \\ c^2=a^2+b^2\implies \pm\sqrt{c^2-a^2}=b\) we're told that the angle is in the 1st Quadrant, in the 1st Quadrant, recall "x" and "y" or "cosine" and "sine" are positive, check your Unit Circle, thus the given cosine function is positive and thus the sine function will also be positive

OpenStudy (jdoe0001):

so once you find what the \(\bf sin(\theta)\) is, then you can find what \(\bf sin(2\theta), cos(2\theta)\) are

OpenStudy (jdoe0001):

recall that \(\bf cos(2\theta)\implies 2cos^2(\theta)-1\qquad \qquad sin(2\theta)=2sin(\theta)cos(\theta)\)

OpenStudy (jdoe0001):

so you're given the cosine , now you'd just need to find the sine function for the angle, so you can plug them in the Double-Angle trig identities

OpenStudy (anonymous):

i cant do i

OpenStudy (jdoe0001):

hmmm can't find \(\bf sin(\theta)\quad ?\)

OpenStudy (anonymous):

:(

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{15}{17}\implies \cfrac{\textit{adjacent side}}{hypotenuse}\implies \cfrac{a}{c}\qquad thus\\ \quad \\ a = 15\qquad c= 17\qquad b=\square ?\qquad \textit{using pythagorean theorem}\\ \quad \\ c^2=a^2+b^2\implies \pm\color{red}{\sqrt{c^2-a^2}=b}\) keep in mind that \(\bf sin(\theta)=\cfrac{b}{c}\)

OpenStudy (anonymous):

8/17

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

sin^2theta - cos^2 theta =1

OpenStudy (jdoe0001):

now you know what the cosine and sine values are thus \(\bf cos(2\theta)\implies 2cos^2(\theta)-1\qquad \qquad sin(2\theta)=2sin(\theta)cos(\theta)\)

OpenStudy (jdoe0001):

hmmm are you after the \(\bf cos^2(\theta) \ or \ cos(2\theta)\quad ?\)

OpenStudy (anonymous):

im rally confused :(

OpenStudy (anonymous):

1-225/289= 289/289-225/289= (289-225)/289=64/289

OpenStudy (jdoe0001):

well... what part confuses you?

OpenStudy (anonymous):

Sho ( your nickname is too long i will just call you that ,you don't mind do you ? ) the equasion is sin^2theta + cos^2 theta =1. You add the values not substract them. From that you learn the value of sin^2theta, from which you know the value of sintheta. Now that you know both sinthetha and costhetha, you use their values while using the formula that jdoe gave you : cos(2θ)⟹2cos^2(θ)−1 sin(2θ)=2sin(θ)cos(θ)

OpenStudy (anonymous):

i cant do it

OpenStudy (anonymous):

what do i have to do after when i got sin 8/17

OpenStudy (jdoe0001):

well, are after \(\bf cos^2(\theta) \ or \ cos(2\theta)\quad ?\)

OpenStudy (anonymous):

i dont know which way i have to start

OpenStudy (anonymous):

You know the values of cosT and sin T. There are said formulas : cos(2θ)=2cos^2(θ)−1 sin(2θ)=2sin(θ)cos(θ) Now you just substitue the functions with their values .

OpenStudy (anonymous):

cos 2 theta =2 cos^2 theta 15/17*2=2(15/17)^2

OpenStudy (anonymous):

No, cos2T=2*cos^2T-1. You forgot to substract the 1.

OpenStudy (anonymous):

30/17=161/289

OpenStudy (anonymous):

cos2T=2*cos^2T-1=2*(15/17)^2-1=239/289 and sin2T=2*cosT*sinT=2*(8/17)*(15/17)=240/289

OpenStudy (anonymous):

cos 2 T = 161/289